_C8H18 + _ O2 -> _CO2 + _H2O
If a gallon of gas produces 6350 grams of CO2 what mass of octane was originally present in one gallon of gas?
1
Expert's answer
2019-03-01T08:15:56-0500

Answer:

2 C8H18 + 25 O2 = 16 CO2 + 18 H2O

n (CO2) = m/M

n (CO2) = 6350/44 = 144.32 mole

n (C8H18) = 144.32*2 / 16 = 18.04 mole

m(C8H18) = 18.84*114.2 = 2151,5 g